OpenStudy (anonymous):

given g(x)=x^3-x find g(-3)

OpenStudy (matt101):

g(-3) = (-3)^3 - (-3) = -27 + 3 = -24

OpenStudy (anonymous):

\[\LARGE {\color{Golden} Hello\: I\: Love\: Math}\]

OpenStudy (anonymous):

hello, i dont

OpenStudy (anonymous):

Sorry man, i am trying to work on Latex Writing:)

OpenStudy (anonymous):

\[\LARGE {\color{Green} YoYoYo}\]

OpenStudy (anonymous):

How come the color wont go for all the words

OpenStudy (anonymous):

its k, but i dont know what that is eitherr

OpenStudy (anonymous):

Its a special way to write math on the computer.

OpenStudy (anonymous):

ohh i see

OpenStudy (anonymous):

\[\LARGE {\color{DarkBlue} Hello\mathbf{YoYo}}\]

OpenStudy (anonymous):

\[\LARGE {\color{DarkBlue}{HelloYoYo}}\]

OpenStudy (anonymous):

Oh i got the color:)

OpenStudy (anonymous):

yess you did

OpenStudy (anonymous):

\[\LARGE {\color{Orange}{Olli\: Olli\: Oxenfree}}\]

OpenStudy (anonymous):

\[\huge \textbf{Hey\, Buddy}\]

OpenStudy (anonymous):

\[\huge \begin{bmatrix} 3 5 4 t \\ & & & & \\ & & & & \\ & & & & \end{bmatrix}\]